OpenStudy (katb):

find the solution to dT-k(T+45)dt=0 that satisfies T(0)=200

OpenStudy (amistre64):

looks separable to me

OpenStudy (amistre64):

\[\frac{1}{T+45}\ dT=-k\ dt\]

OpenStudy (amistre64):

but without the -

OpenStudy (amistre64):

drop an integration on it to get: \[ln(T+45)=kt + C\]

OpenStudy (katb):

how do I continue?

OpenStudy (amistre64):

"e" it up and isolate for T

OpenStudy (amistre64):

im not sure how well that k is going to play tho; but since t = 0 we can assume its not really a factor in the overall scheme

OpenStudy (katb):

and then how do i finish and find T(0)=200

OpenStudy (amistre64):

well, after you isolate the T; you make t=0 and solve for C

OpenStudy (amistre64):

t=0 and T=200 that is

OpenStudy (amistre64):

you might be able to play it as is; but lns tend to get persnickety